3. If f(x) = 2x + 4 and g(x) = - 3. Given that h(x) = g(x) – f(x). Find h(x).
1
Expert's answer
2014-05-15T09:30:05-0400
Substituting we obtain
h(x) = g(x) – f(x) = - 3 - (2x + 4) = - 3 - 2x - 4 = -2x - 7.